# Drivetrain simulation using variational autoencoders

### v1

@@ 5 Conclusion
In this study, we showed the potential of generative artificial intelligence (AI) for drivetrain simulation. Conventional supervised models map drivetrain behavior based on input settings, but our work introduces a self-supervised approach using variational autoencoders (VAEs). We demonstrated how VAEs encode jerk signal information into a latent space, enabling the generation of new meaningful jerk scenarios through sampling. Extending this concept, we implemented two conditional VAEs to address simulation tasks that depend on a specific driver torque demand. In particular, we could conclude the following:
@@
  • When measuring the absolute error between original and generated spectrograms, localized errors of up to 25% are observed. Initially, this might suggest poor reconstruction, but a closer look reveals that the original spectrograms appear noisy, whereas the generated ones are smoother. This suggests that the VAE filters out noise and captures the core features. After applying a Gaussian filter to smooth the original spectrograms, the absolute error decreased, confirming this hypothesis. However, training the VAE with smoothed spectrograms generated signals with unexpected frequency peaks, deviating from the dominant 1 Hz and 10 Hz frequencies typically seen. Thus, although smoothing reduces error, it does not improve the VAE’s generative capability. The VAE’s ability to generate meaningful signals from noisy data suggests that it effectively captures the essential features, showing resilience to noise. No significant differences in KL divergence were observed during training when using the original or smoothed spectrograms.
@@
To remove the unwanted effect of jerk, one can calibrate the engine maps. For this purpose, one can build a physics-based hardware simulator that simulates the jerk and acceleration given a torque signal as input. Theoretically, the simulator could create signals close to the real world, but complex physical interactions affect the accuracy of the simulations and cause them to diverge from real measurements. One can build a hybrid hardware simulator by using a correction model [52, 53, 54]. The correction model is used to bring the simulated signal close to reality. To do so, a neural network trained on [simulated acceleration; measured acceleration] pairs can be used, as we choose to use the acceleration signal for calibration purposes.

### v2

@@ 5 Conclusion
In this study, we showed the potential of deep generative models, specifically VAEs and their conditional variants, for drivetrain simulation. Conventional supervised models map drivetrain behavior based on input settings, but our work introduces a self-supervised approach using VAEs. We demonstrated how VAEs encode jerk signal information into a latent space, enabling the generation of new meaningful jerk scenarios through sampling. Extending this concept, we implemented two conditional VAEs to address simulation tasks that depend on a specific driver torque demand. In particular, we could conclude the following:
@@
  • When measuring the absolute error between original and generated spectrograms, localized errors of up to 25% are observed. Initially, this might suggest poor reconstruction, but a closer look reveals that the original spectrograms appear noisy, whereas the generated ones are smoother. This suggests that the VAE filters out noise and captures the core features. After applying a Gaussian filter to smooth the original spectrograms in an auxiliary analysis, the absolute error between the VAE’s reconstruction and these smoothed spectrograms decreased, confirming the VAE’s tendency to produce smooth outputs. However, when we trained a VAE model using these artificially smoothed spectrograms as input, the subsequently generated signals sometimes exhibited unexpected frequency peaks, deviating from the dominant 0-2 Hz and 8-12 Hz frequencies typically seen in the original data. Thus, this exploratory experiment suggested that although pre-smoothing the input data can reduce certain reconstruction error metrics, it does not necessarily improve the VAE’s generative capability for physically plausible signals from the original noisy distribution and might even hinder it. The VAE’s ability to generate meaningful signals from noisy data suggests that it effectively captures the essential features, showing resilience to noise. No significant differences in KL divergence were observed during training when using the original or smoothed spectrograms.
@@
To remove the unwanted effect of jerk, one can calibrate the engine maps by building a physics-based hardware simulator that simulates the jerk and acceleration given a torque signal as input. Theoretically, the simulator could create signals close to the real world, but complex physical interactions affect the accuracy of the simulations and cause them to diverge from real measurements. As a consequence, one can build a hybrid hardware simulator by adding a correction model [53, 54, 55]. The correction model is used to bring the simulated signal close to reality by learning complex physical interactions from the data. To do so, a neural network trained on [simulated acceleration; measured acceleration] pairs can be used, as we choose to use the acceleration signal for calibration purposes.